OpenStudy (kamille):

I need to factorise this: \[\frac{ (5n-2)! }{ (5n-4)! }\] can someone help me?

OpenStudy (asnaseer):

do you know the definition of x! ?

OpenStudy (kamille):

yes, i do

OpenStudy (kamille):

maybe:

OpenStudy (asnaseer):

use that in the numerator and you should see what to do...

OpenStudy (kamille):

\[\frac{ (5n-2)! }{ (5n-4)(5n-3)(5n-2)! }\]

OpenStudy (kamille):

but i am not sure

OpenStudy (asnaseer):

not quite

OpenStudy (asnaseer):

x! = x(x-1)(x-2)...(1)

OpenStudy (asnaseer):

so (5n-2)! = (5n-2)( (5n-2) - 1)( (5n-2) - 2)...(1)

OpenStudy (kamille):

oh, i see. thank you a lot!

OpenStudy (asnaseer):

yw :)